OpenStudy (anonymous):

[7.03] Choose the correct simplification of 9y^0.

OpenStudy (anonymous):

9y^0 = 9 , am i correct ?

OpenStudy (anonymous):

would it be 0

OpenStudy (anonymous):

nm i got it

OpenStudy (anonymous):

uhmm .. i dont know I doubt it .. @Hero plis help

OpenStudy (anonymous):

@greysica your're correct since anything raised to 0 is 1

OpenStudy (anonymous):

yea it is 1

OpenStudy (anonymous):

BY THING RAISER TO POWER ZERO IS ONE so 1*9 will be???

OpenStudy (anonymous):

not 0

OpenStudy (anonymous):

or 9

OpenStudy (anonymous):

ANY*

OpenStudy (anonymous):

@savannajadey its 9*1 =???

OpenStudy (anonymous):

1

OpenStudy (anonymous):

nine times one = nine

OpenStudy (anonymous):

IT IS 9 :p

OpenStudy (anonymous):

thanks you @gitahimart @Hero i was doubt

OpenStudy (anonymous):

nooooooooooo anything with an exponent og 0 is 1

OpenStudy (anonymous):

so one multiplied by 9 is??

OpenStudy (anonymous):

ow

OpenStudy (anonymous):

it is 9 * y^0

OpenStudy (anonymous):

yes y^0 = 1 but we have 9y^0 so 9*1 :)

OpenStudy (anonymous):

y^0 =1

OpenStudy (anonymous):

true

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

if it was (9y)^0 then the answer would be 1, but your question is 9y^0 = 9* y^0 = 9*1 = 9